The wild thing has a mass of 500 kg with no passengers, at the top of the biggest hill the ride is 63m in the air and moving at 1m/s. Find the gravitational potential energy, kinetic energy and total energy at the top of the biggest hill.
1
Expert's answer
2019-02-17T09:36:07-0500

Gravitational potential energy at the top of the biggest hill:

E_p=mgh=500∙9.8∙63=308700 (J)

Kinetic energy at the top of the biggest hill:

E_k=(mϑ^2)/2=(500∙1^2)/2=250 (J)

Total energy:

E=E_p+E_k=308700+250=308950 (J)